the present age of john = the square of the age of his son,last year the age of john was 8 times the age of his son,what are their respective ages.
Answer by ankor@dixie-net.com(22740)

let j = John's present age
Let s = son's present age
:
Write an equation for each statement:
:
"the present age of john = the square of the age of his son,"
j = s^2
:
"last year the age of john was 8 times the age of his son,"
j - 1 = 8(s-1)
j - 1 = 8s - 8
j = 8s - 8 + 1
j = 8s - 7
Replace with s^2
s^2 = 8s - 7
arrange as a quadratic equation
s^2 - 8s + 7 = 0
Factors to
(s-7)(s-1) = 0
Two solutions
s = 1, not reasonable since a year ago he would be 0
s = 7 yrs is the son's age
then
j = 7^2
j = 49 is John's age
:
:
See if that checks out in the statement:
"last year the age of john was 8 times the age of his son,"
49 - 1 = 8(7-1)
48 = 8(6)